# Disentangling Boltzmann brains, the time-asymmetry of memory, and the second law

### v3

@@ I Introduction
As we show, this simple observation, formulated in a fully rigorous manner, disentangles all the ambiguities and circular reasoning that are rife in the literature on Boltzmann brains, and more generally in the literature on the past hypothesis (PH), and even in the literature on the second law. In particular, our analysis shows that the PH and the BB hypothesis are, formally speaking identical, in that both only condition on a subset of our current data, differing only in which such data they condition on. We also show that arguments about “unstable reasoning” prove nothing concerning the validity of the BB hypothesis, only about one particular argument that has been offered for that validity. We also formalize some of the properties of all circular reasoning, illustrating that they apply to common arguments for the second law.
@@
We begin in the next section by summarizing one of the more nuanced, semi-formal counter-arguments that have been made to the BB hypothesis. A central concern with the counter-argument presented there is how we can reason from current data to conclusions about how experiments were initialized in the past, and from there to infer laws of the universe. The key conundrum we highlight is that this inference process itself relies on those laws being inferred, and so is circular. (Arguments in the literature that reach similar conclusions have characterized the standard argument for the BB hypothesis as “ unstable reasoning” Myrvold (2016); Carroll (2016) for some work related to this counter-argument.)
@@
Our third contribution is to use this formalization of the entropy conjecture to clarify what it cannot tell us. In contrast, our fourth contribution is to clarify what the entropy conjecture does in fact tell us concerning the BB hypothesis specifically. Next, our fifth contribution is to use our formalization of the entropy conjecture to analyze what arguments concerning unstable reasoning can and cannot tell us about the BB hypothesis. After this we present our sixth contribution, which is a variant of the BB hypothesis we call the “1000 CE hypothesis”. As we show, formally speaking, the 1000 CE hypothesis lies exactly between the standard BB hypothesis and the PH. This illustrates the underlying identity between standard BB hypothesis and the PH.
@@ III Relation between the entropy conjecture, Boltzmann Brains, and the second law
The discussion in the preceding sections may leave a sense of confusion. Many parts of the arguments recounted above involving the BB hypothesis might seem to involve circular reasoning. Such reasoning can only (!) establish that our priors are inconsistent. However, the fact that an argument engages in circular reasoning doesn’t tell us anything about how valid its conclusions are, just that the priors it relies on don’t themselves establish those conclusions. (See Appendix B for a Bayesian definition “circular reasoning”, and an analysis of what it does and does not imply.)
Another possible concern is that the arguments are only semi-formal, relying on “common experience”. However, in any investigation concerning foundational issues in thermodynamics, (like whether BBs accord with our present data records), if one does not use fully formal mathematical reasoning it is extremely easy to subtly assume what one is in fact trying to prove. For example, that mistake is made in very many of the papers in the literature trying to prove the second law. Specifically, the analyses in many of these papers implicitly privilege the initial condition of a physical system (e.g., how it was set up) over its final condition (e.g., how it finishes). That asymmetry in how it specifies boundary conditions is ultimately the basis for the “derivation” of the second law. However, privileging the initial conditions in turn arises from the psychological arrow of time – which it is commonly believed to be a consequence of the asymmetry of the second law. So such a supposed derivation involves circular reasoning. The same mistake of circular reasoning has also been made in many earlier investigations of the BB hypothesis.
@@ III.1 Using probability distributions to reason about Newton’s laws
We label the first such observational data set we are interested in as ${\cal{D}}_{1}$. ${\cal{D}}_{1}$ consists of current scientific records which lead to the conclusion that our universe obeys Newtonian mechanics (assuming those records do in fact provide the result of the associated experiments, which are in our past). In other words, they are the current state of some physical “memory system” that is statistically correlated with the state of some experimental outcomes in the past Wolpert and Kinney (2024), where those experiments involved Newton’s laws. Stated more formally, if ${\cal{D}}_{1}$ — the current state of an associated physical memory system — does in fact accurately provide the results of experiments conducted in our past, then $P(\mbox{Newton's laws hold}\,|\,\mbox{observational data }{\cal{D}}_{1})$ would be high. (In this paper we will not need to precisely quantify what it means for some probability to be “high”.) Below we will refer to this situation as ${\cal{D}}_{1}$ being “reliable”.
@@
To get past this roadblock we need to in essence assume that ${\cal{D}}_{1}$ does indeed accurately provide the results of past experiments. More precisely, we need to adopt a prior distribution over $\Gamma$ that says that we can indeed use the records ${\cal{D}}_{1}$, which are the current state of an associated memory system, to infer that with high probability the results of those earlier experiments, which in turn imply Newton’s laws. While these details don’t concern us here, it is important to note that can make this reasoning concerning prior distributions and our data fully formal. To do so, in the terminology of Wolpert and Kipper (2024) we say that ${\cal{D}}_{1}$ is reliable if it is the current state of a memory system, and there is high restricted mutual information between ${\cal{D}}_{1}$ and the results of those particular experiments in our past which are related to Newton’s laws. [^3] Formally, “reliability” is an assumption about the nature of the physical process coupling the state of the memory system that contains the value ${\cal{D}}_{1}$ and the state of (some variables in) the universe in the past. ${\cal{D}}_{1}$ is, strictly speaking, a value of the random variable given by the current state of the memory system — and reliability is a property of the latter. To make this more precise, consider the following change of notation. Let ${\cal{D}}_{1}$ be that random variable, taking values in a set $D_{1}$, with a precise sample of the associated distribution written as $d_{1}\in D_{1}$. Suppose there is a set of possible values of $S^{{\cal{D}}_{1}}\subseteq D_{1}$ such that if $d_{1}\in S^{D_{1}}$, then due to prior assumptions, we know that the posterior $P(\mbox{Newton's laws hold}\,|\,d_{1}\in S^{D_{1}})$ is high. (The precise value of that posterior of course can be higher or lower, depending on what precisely $d_{1}$ is.) See appendix on unstable reasoning.
@@ III.2 The entropy conjecture
Unfortunately, this precise failure to formulate arguments concerning the dynamics of the universe’s entropy as a stochastic process has led to widespread failure to fully grasp the mathematics of the entropy conjecture, and how to use it to reason formally about the second law, BBs, etc.. To undertake such reasoning one must formulate the entropy conjecture in terms of stochastic processes.
@@
As an aside, an important property of Boltzmann processes is that their marginal distribution must be the same at all times $t$ i.e., $P(x_{t})$ must be independent of $t$. (This follow from the fact that the kernel is time-translation invariant.) It immediately follow that that marginal distribution equals the fixed point distribution of the kernel Scharnhorst et al. (2024).
@@ III.3 The Boltzmann Brain hypothesis and the entropy conjecture
In light of this clarification, is there any way we might infer the BB hypothesis from the assumption only (that ${\cal{D}}_{1}$ is true and therefore) the entropy conjecture holds, as the standard arguments for the BB hypothesis in the literature claim to do? To answer this question, first define ${\cal{D}}_{2}$ to be some set of scientific records which have the property that if the entropy conjecture holds (i.e., if the first supposition of the BB hypothesis holds), and ${\cal{D}}_{2}$ is reliable, then the other two suppositions of the BB hypothesis hold with high posterior probability, conditioned on both ${\cal{D}}_{1}$ and ${\cal{D}}_{2}$. We will use the term “standard argument” for the BB hypothesis to mean the assumption that there is such a ${\cal{D}}_{2}$, and so suppositions (ii) and (iii) of the BB hypothesis hold.
@@
As an aside, note that before the advent of observational data indicating the existence of the Big Bang in our past, we did not have the data set ${\cal{D}}_{3}(2)$, only ${\cal{D}}_{3}(1)$. In those times before Hubble, arguably scientists should have concluded (absent other arguments) that we were Boltzmann brains. Alas, scientists of the time did not have the courage of their convictions.
@@ III.4 The Past Hypothesis and the Second Law
So a prior that ${\cal{D}}_{1}$ and ${\cal{D}}_{3}(2)$ are reliable, together with the assumption that item (iii) of the PH holds, and no other data, would imply that the PH holds, and that therefore the second law holds. Note that the second law would in turn mean that entropy decreases the further back we go into our past from the present. However, as mentioned above, the BB hypothesis would imply that entropy increases going backward in time from the present. So the PH would both imply the second law and rule out the BB hypothesis.
@@ III.5 What the entropy conjecture does not imply
So the second law is consistent with the priors that imply it — but only by virtue of the fact that deriving it relies on circular reasoning. We can adopt priors that ${\cal{D}}_{1}$ and ${\cal{D}}_{3}$ are reliable, or that the second law holds (or even both). But we cannot somehow avoid making any prior assumptions at all, in an attempt to bootstrap ourselves into placing high probability on having ${\cal{D}}_{1}$ or ${\cal{D}}_{3}$ be reliable and that the second law of thermodynamics holds. (See Section B.1 for a Bayesian formalization of circular reasoning, and associated derivation of what circular reasoning does and does not imply.)
Suppose that we misunderstood the foregoing, and thought that somehow the reliability of ${\cal{D}}_{1}$ not only implies the entropy conjecture, but also implies:
@@
    If the second law holds (both at the present and at times in our past), then it follows that ${\cal{D}}_{1}$ is reliable. (In other words, there is a type-3 memory system that would rely on the second law in order to establish that ${\cal{D}}_{1}$ is reliable.)
@@
However, this derivation of the BB hypothesis would mean that the second law does not hold. That would contradict the second supposition (b) of this very derivation of the BB hypothesis. Note that this is more than using circular reasoning to try to establish some hypothesis or other. Rather it is using reasoning that contradicts itself to try to establish a particular hypothesis. It is this seeming self-contradiction that we presented in Section II, and is related to unstable reasoning.
@@
IV Discussion

### v4

@@ I Introduction
As we show, this simple observation, formulated in a fully rigorous manner, disentangles all the ambiguities and circular reasoning that are rife in the literature on Boltzmann brains, and more generally in the literature on the past hypothesis (PH), and even in the literature on the second law. In particular, our analysis shows that the PH and the BB hypothesis are, formally speaking identical, in that both only condition on a subset of our current data, differing only in which such data they condition on. We also show that arguments about “unstable reasoning” do not in fact prove anything concerning the validity of the BB hypothesis, only about one particular argument that has been offered for that validity. We also formalize some of the properties of all circular reasoning, illustrating that they apply to common arguments for the second law.
@@
We begin in Section II by summarizing one of the more nuanced, semi-formal counter-arguments that can be made to the BB hypothesis. A central concern with the counter-argument presented there is how we can reason from current data to conclusions about how experiments were initialized in the past, and from there to infer laws of the universe. The key conundrum we highlight is that this inference process itself relies on those laws being inferred, and so is circular. (Arguments in the literature that reach similar conclusions have characterized the standard argument for the BB hypothesis as “unstable reasoning”. See Myrvold (2016); Carroll (2016) for some work related to this counter-argument.)
@@
Our third contribution is to use this formalization of the entropy conjecture to clarify what it cannot tell us. In contrast, our fourth contribution is to clarify what the entropy conjecture does in fact tell us concerning the BB hypothesis specifically. Next, our fifth contribution is to use our formalization of the entropy conjecture to formalize the reasoning in Section II, and to analyze what arguments concerning unstable reasoning can and cannot tell us about the BB hypothesis. After this we present our sixth contribution, which is a variant of the BB hypothesis we call the “1000 CE hypothesis”. As we show, formally speaking, the 1000 CE hypothesis lies exactly between the standard BB hypothesis and the PH. This illustrates the underlying identity between the standard BB hypothesis and the PH.
@@ III Relation between the entropy conjecture, Boltzmann Brains, and the second law
The discussion in the preceding sections may leave a sense of confusion. Many parts of the arguments recounted above involving the BB hypothesis might seem to involve circular reasoning. Such reasoning can only (!) establish that the priors it relies on are inconsistent. However, the fact that an argument engages in circular reasoning doesn’t tell us anything about how valid its conclusions are, just that the priors it relies on don’t themselves establish those conclusions. (See Appendix B for a Bayesian definition “circular reasoning”, and an analysis of what it does and does not imply.)
Another possible concern is that the arguments are only semi-formal, relying on “common experience”. However, in any investigation concerning foundational issues in thermodynamics, (like whether BBs accord with our present data records), if one does not use fully formal mathematical reasoning it is extremely easy to subtly assume what one is in fact trying to prove. For example, that mistake is made in very many of the papers in the literature trying to prove the second law. Specifically, the analyses in many of these papers implicitly privilege the initial condition of a physical system (e.g., how it was set up) over its final condition (e.g., how it finishes). That asymmetry in boundary conditions assumed by those arguments is, ultimately, the basis for the “derivation” of the second law. However, our privileging the initial conditions in turn arises from the psychological arrow of time – which it is commonly believed to be a consequence of the asymmetry of the second law. So such a supposed derivation involves circular reasoning. The same mistake of circular reasoning has also been made in many earlier investigations of the BB hypothesis.
@@ III.1 Using probability distributions to reason about Newton’s laws
We label the first such observational data set we are interested in as ${\cal{D}}_{1}$. ${\cal{D}}_{1}$ consists of current scientific records which lead to the conclusion that our universe obeys Newtonian mechanics (assuming those records do in fact provide the result of the associated experiments, which are in our past). In other words, they are the current state of some physical “memory system” that we suppose (!) is statistically correlated with the state of some experimental outcomes in the past Wolpert and Kinney (2024), where those experiments involved Newton’s laws. Stated more formally, if ${\cal{D}}_{1}$ — the current state of an associated physical memory system — does in fact accurately provide the results of experiments conducted in our past, then $P(\mbox{Newton's laws hold}\,|\,\mbox{observational data }{\cal{D}}_{1})$ would be high. (In this paper we will not need to precisely quantify what it means for some probability to be “high”.) Below we will refer to this situation as ${\cal{D}}_{1}$ being “reliable”.
@@
To get past this roadblock we need to in essence assume that ${\cal{D}}_{1}$ does indeed accurately provide the results of past experiments. More precisely, we need to adopt a prior distribution over $\Gamma$ that says that we can indeed use the records ${\cal{D}}_{1}$, which are the current state of an associated memory system, to infer that with high probability the results of those earlier experiments, which in turn imply Newton’s laws. While these details don’t concern us here, it is important to note that can make this reasoning concerning prior distributions and our data fully formal. To do so, in the terminology of Wolpert and Kipper (2024) we say that ${\cal{D}}_{1}$ is reliable if it is the current state of a memory system, and there is high restricted mutual information between ${\cal{D}}_{1}$ and the results of those particular experiments in our past which are related to Newton’s laws. [^3] Formally, “reliability” is an assumption about the nature of the physical process coupling the state of the memory system that contains the value ${\cal{D}}_{1}$ and the state of (some variables in) the universe in the past. ${\cal{D}}_{1}$ is, strictly speaking, a value of the random variable given by the current state of the memory system, and reliability is a property of the coupling.
It is important to emphasize that every paper concerning the second law, the past hypothesis, or Boltzmann brains, has assumed the laws of physics hold. Many of those papers go on to make additional assumptions as well. Here we are maximally conservative, not making any extra assumptions at all.
@@ III.2 The entropy conjecture
Unfortunately, this precise failure to formulate arguments concerning the dynamics of the universe’s entropy as a stochastic process has led to widespread failure to fully grasp the mathematics of the entropy conjecture, and how to use it to reason formally about the second law, BBs, etc. To undertake such reasoning one must formulate the entropy conjecture in terms of stochastic processes.
@@
As an aside, an important property of Boltzmann processes is that their marginal distribution must be the same at all times $t$ i.e., $P(x_{t})$ must be independent of $t$. (This follow from the fact that the kernel is time-translation invariant.) It immediately follows that that marginal distribution equals the fixed point distribution of the kernel Scharnhorst et al. (2024).
@@ III.3 The Boltzmann Brain hypothesis and the entropy conjecture
In light of this clarification, is there any way we might infer the BB hypothesis from the assumption only that (${\cal{D}}_{1}$ is true and therefore) the entropy conjecture holds, as the standard arguments for the BB hypothesis in the literature claim to do? To answer this question, first define ${\cal{D}}_{2}$ to be some set of scientific records which have the property that if the entropy conjecture holds (i.e., if the first supposition of the BB hypothesis holds), and ${\cal{D}}_{2}$ is reliable, then the other two suppositions of the BB hypothesis hold with high posterior probability, conditioned on both ${\cal{D}}_{1}$ and ${\cal{D}}_{2}$. We will use the term “standard argument” for the BB hypothesis to mean the assumption that there is such a ${\cal{D}}_{2}$, and so suppositions (ii) and (iii) of the BB hypothesis hold.
@@
As an aside, suppose that before the advent of observational data indicating the existence of the Big Bang in our past, we did not have the data set ${\cal{D}}_{3}(2)$, only ${\cal{D}}_{3}(1)$. Suppose that in those times before Hubble, scientists had our current understanding of the Big Bang, the PH, why memory is temporally asymmetric, etc., Arguably, if that had been the case, then those scientists should have concluded (absent other arguments) that we were Boltzmann brains. See Section IV.
@@ III.4 The Past Hypothesis and the Second Law
[…]
@@ III.5 What the entropy conjecture does not imply
So the second law is consistent with the priors that imply it — but only by virtue of the fact that we can adopt priors that ${\cal{D}}_{1}$ and ${\cal{D}}_{3}$ are reliable, or that the second law holds (or even both). But we cannot somehow avoid making any prior assumptions at all, in an attempt to bootstrap ourselves into placing high probability on having ${\cal{D}}_{1}$ or ${\cal{D}}_{3}$ be reliable and that the second law of thermodynamics holds. (See Section B.1 for a Bayesian formalization of circular reasoning, and associated derivation of what circular reasoning does and does not imply.)
Suppose that we misunderstood the foregoing, and thought that somehow if ${\cal{D}}_{1}$ were reliable, that would not only imply the entropy conjecture, but also imply:
@@
    If the second law holds (both at the present and at times in our past), then it follows that in fact ${\cal{D}}_{1}$ is reliable. (In other words, there is a type-3 memory system that would rely on the second law in order to establish that ${\cal{D}}_{1}$ is reliable.)
@@
However, this derivation of the BB hypothesis would mean that the second law does not hold. That would contradict the second supposition (b) of this very derivation of the BB hypothesis. Note that this is more than using circular reasoning to try to establish some hypothesis or other. Rather it is using reasoning that contradicts itself to try to establish a particular hypothesis. This seeming self-contradiction is a version of unstable reasoning; see Appendix B.
@@
In addition though, it is interesting to note that nobody has ever invoked the second law as in (b) to justify the reliability of any data concerning the past they wish to use, be that data ${\cal D}_{1}$ or anything else. Rather, we now understand that all memory systems that allow us to treat the current state of a memory as providing information concerning the past must, ultimately, rely on the second law. [^6] Indeed, when the memory system in question is the human brain, we have no idea of the details of how the second law establishes the validity of memory. We just believe that this must be the case. But before this understanding, say a century ago, there would have been no reason to suppose that the second law was in any way required for us to (have faith that we can) use our experimental data as though it provides accurate information about the past state of our experimental apparatuses. There would not have been any reason to suppose that the second law is itself in any way needed to derive the second law itself. And no current textbook uses the second law that way. “Unstable reasoning” never explicitly arises in any of the arguments for the BB hypothesis one can find in the literature.
@@
IV How and why intuition leads us astray
Many of the mathematical arguments and derivations presented above defy to common sense and intuition. In part, one might dismiss this issue. After all, since the advent of quantum physics and the general theory of relativity, physicists have become accustomed to following math wherever it may lead, even when it results in counterintuitive conclusions.
Nonetheless, it is worth teasing apart the precise reason why our result are counterintuitive, to allay associated fears. Some instances of specific intuitions being defied were discussed above, in both the introduction and Section III. Since these are such subtle points however, It’s worth describing some of the other strong intuitions which are at odds with our formal results, why those intuitions are so strong, and just how it is that they can be so misleading.
First, most obviously, almost everyone feels, very deeply, that their memories are accurate, that they correctly depict some information about a real external world and its states in the recent past. To understand how this deep-felt intuition could be completely wrong, it’s necessary to proceed through several steps, to understand the source of this intuition:
  1. 1.
    To begin, note that human memory is, formally speaking, a physical system (namely, our brains) whose current state provides statistical information about the likely state of some external physical system, at a different time (namely, information about the external world that we believe we remember). Or at least, we presume that to be the case.
    Formalizing memory this way, we immediately notice a mystery: why do we have so much confidence that the current state of our brains provides such accurate information about the past state of the external world, but not about the future state of the external world? Phrased differently, why do we believe our innate ability at retrodiction is so much greater than our ability at prediction?
    Presumably, our belief in the accuracy of our memories of the past does not have anything to do with our knowledge of the laws of physics. Training in theoretical thermodynamics is certainly not required for people to believe their memories are accurate (!). Rather, this belief seems to be built into the functioning of our brains, into how our neocortices are wired, so to speak. Why are our brains constructed this way, to believe our memories of our past, solely?
  2. 2.
    To address that mystery requires us to provide a physical model of memory. However, human memory involves some extraordinarily complicated processes in the human brain, processes we know precious little about. [^10] We know for example that the hippocampus plays a crucial role, and that elaborations of Hebbian learning, like post-synaptic plasticity, play a major role at the level of individual neurons. But even this much is understood only in generalities.
    Accordingly, the great majority of investigations of the human memory system have not considered it directly. Rather they have considered far simpler examples of memory systems that we can investigate in detail, which provide non-biological instances of memory Wolpert (1992); Wolpert and Kipper (2024); Rovelli (2014, 2022); Reichenbach (1991); Davies (1977). A common example of such a memory system is a the surface of the moon. The surface of the moon is the memory system, akin to the human brain, whose current state (namely, the craters on the moon’s surface) provides information about the state of physical systems external to it in the past (namely, that there were meteors on a collision course with the moon in the past). Another common example is a line of footprints on an otherwise smooth beach. The beach is the memory system, akin to the human brain, whose current state (namely, the footprints on it) provides information about the state of physical systems external to it in the past (namely, that there was a person walking across that beach).
  3. 3.
    What do we learn from these examples of simple memory systems concerning the asymmetry of our memory? Do the common features of these examples of memory systems help us understand why we have so much more confidence in retrodiction rather than prediction?
    In fact, these examples do have one striking feature in common: all of them seem to rely on the second law of thermodynamics. The temporal asymmetry of these memory systems, and by extension of our human brains, seems to reflect nothing more than the simple fact that entropy increases in time. Indeed, the precise reason we are so very sure about some of our memories seems to reflect the ironclad nature of the second law.
    In other words, the reason we seem to have so much confidence in our memory ultimately resides in the fact that those memories rely on the second law.
  4. 4.
    This line of reasoning would seem to mean that we should have complete faith in our memories, given that the second law cannot be violated. Unfortunately though, this line of reasoning begs a question: how it is that the physical universe, whose (relevant) microscopic laws are all time-reversal symmetric, could give rise to the asymmetry of the second law? The arguments supporting the second law offered by Boltzmann and those who followed him were all based on the microscopic laws of physics, and therefore they had to be time-inversion symmetric. Formally they all state that f at a certain time $t$ entropy is low (or a similar quantity is, like Boltzmann’s H function), then entropy increases in time going both forward from $t$ and backward from $t$ Davies (1977). This implies that entropy should increase going both forward from the present as well as going backward from the present.
    In other words, our analysis of why we believe our memories to be accurate has reduced the mystery of human memory to the infamous paradox ascribed to Loschmidt.
  5. 5.
    What would this time-symmetric reasoning of Loschmidt (and others) mean physically? What would it mean, physically, for entropy to increase going into our past, despite our apparent memories of the past, which would seem to rely on entropy increasing in time, not decreasing? Among other things, it would have to mean that we have no reason to believe our memories are accurate. It means that they do not tell us anything about the state of the external world in the past. They only appear to give us that information.
    In short, our intuition concerning our memories leads us greatly astray. In quantum mechanics, or relativity, that which seems “obviously” true is deeply problematized, either by the results of experiment (quantum mechanics) and / or pure mathematical reasoning (general relativity). The same is true of thermodynamics, evidently.
    Just like when we analyze event horizons or the quantum Zeno effect, when when we analyze the second law, memory and Boltzmann brains, we are almost guaranteed to go astray if we don’t rely solely on mathematical reasoning to guide us, consigning intuition to a purely secondary role.
Reasonable as the arguments just presented might be, in the abstract, how, concretely, can they hold? How could we have all of our human memories concerning the past be fallacious? How could entropy increase into our past rather than decrease, as required by the time-symmetric nature of all derivations of the second law that are consistent with the microscopic laws of physics? How could it be that our memories are wrong?
Such flaws in our memory would require some exquisite fine-tuning, that all the neurons in our brains happen to be in the state corresponding to particular memories, when in fact nothing of the sort is true. Amazingly though, standard arguments of statistical physics tell us that it is almost infinitely more likely for this to be the case, rather than for entropy to continue to decrease into our past, as demanded by the second law.
This conclusion is precisely the Boltzmann brain hypothesis. As we show above, it too has subtleties that can only be addressed through careful mathematical reasoning.
V Discussion